OpenStudy (anonymous):

(d/4)+4=(d/5)-1/4 what does d equal

OpenStudy (anonymous):

(d/4)+4=(d/5)-1/4 (d/4)+4-4=(d/5)-1/4-4 (d/4)-d/5=(d/5)-1/4-4-d/5 (d/4)-(d/5)=-17/4 (5d-4d)/20=-17/4 (5d-4d)/20*20=-17/4*20 (5d-4d)=-17/4*20=-85 d=-85

OpenStudy (anonymous):

thx